General Physics

A bomber plane moves horizontally with a speed of 500 m/s and a bomb released from it, strikes the ground in 10 sec. angle at which it strikes the ground will be (g=10m/s)

1. tan-1 (1/5)

2. tan-1 (1/2)

3. tan-1 (1)

4. tan-1 (5)

required- answer and solution

cute girls
13 Years agoGrade

1 Answer

sanatan sharma
13 Years ago

 the bomb is dropped  , therefore its horizontal velocity = 500 

and vertical velocity = 0

in 10 sec 

 v = u + at 

   = 0 + 10x10

    = 100 m/s

horizontal velocity remains unchanged  bcoz there''s no accn  in the horizontal 

therefore 

tan Θ =  ( vertical velocity )/ (horizontal velocity )  = 1/5

 

PLEASE LIKE BELOW ON APPROVE IF YOU LIKE THE SOLN